# Gravitational Wave Signal Extraction Against Non-Stationary Instrumental Noises with Deep Neural Network

### v1

@@ Abstract
Sapce-borne gravitational wave antennas, such as LISA and LISA-like mission (Taiji and Tianqin), will offer novel perspectives for exploring our Universe while introduce new challenges, especially in data analysis. Aside from the known challenges like high parameter space dimension, superposition of large number of signals and etc., gravitational wave detections in space would be more seriously affected by anomalies or non-stationarities in the science measurements. Considering the three types of foreseeable non-stationarities including data gaps, transients (glitches), and time-varying noise auto-correlations, which may come from routine maintenance or unexpected disturbances during science operations, we developed a deep learning model for accurate signal extractions confronted with such anomalous scenarios. Our model exhibits the same performance as the current state-of-the-art models do for the ideal and anomaly free scenario, while shows remarkable adaptability in extractions of coalescing massive black hole binary signal against all three types of non-stationarities and even their mixtures. This also provide new explorations into the robustness studies of deep learning models for data processing in space-borne gravitational wave missions.
@@ I INTRODUCTION
Since the first landmark event GW150914 captured by Adv-LIGO [1], today nearly a hundred Gravitational Wave (GW) signals from mergers of stellar mass compact binaries had been observed by the LIGO-VRIGO collaboration, which lead to numerous accomplishments in astrophysics and fundamental physics [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. However, limited by seismic noises, ground-based GW detectors are typically sensitive to signals at frequencies higher than 10 Hz [15, 16, 17]. To enclose the exciting sources of larger and heavier astrophysical systems, one needs to explore the low frequency band of the GW spectrum with detectors of much longer baselines. The first generation space-borne antennas, including LISA (Laser Interferometer Space Antenna) [18, 19] and the LISA-like missions Taiji [20, 21] and Tianqin [22], would be launched in the 2030s and cover the mHz band. Offering novel perspectives for exploring the Universe, such space-borne antennas will also introduce new challenges [19], especially in data analysis [23, 24, 25].
@@
Space antennas are supposed to response to a large amount of superimposed GW signals, including coalescing (super) Massive Black Hole Binaries (MBHB), extreme-mass-ratio-inspirals, galactic compact binaries, evolving topological defects, primordial GW background, and also un-modeled sources. To gain considerable Signal-to-Noise Ratios (SNR) against the overwhelming noises, one generally needs to accumulate data for rather long periods, even for years, which largely increases the computational costs in data processing. The high parameter space dimension of each type of sources adds more complexities and challenges. Many works have been done to address those issues [26, 27, 28]. As lessons learned from the LIGO-VIRGO observatories [29] and precedent missions like GRACE/GFO [30], LISA PathFinder (LPF) [31] and also Taiji-1 [32], new issues in data analysis for space-borne GW detections begin to draw more attentions [33, 34, 35, 36, 25]. As mentioned, to precisely measure the related parameters and infer the physical properties of the sources, continues measurements without disruptions in the data are normally demanded, which however imposes hard challenges on the long-term stability and robustness of both the payloads and satellite platforms. GW detection in space would be more seriously affected by anomalies or non-stationarities in the science measurements, including instrument transients or noise bursts, data gaps, slowly varying noise auto-correlations, trends and etc, as the detailed knowledge of instrumental noises is crucial to GW signals extraction and parameter estimations. According to the designs of the LISA and Taiji missions, the foreseeable anomalies or non-stationarities in science measurements may come from the routine maintenance of the satellites or the onboard payloads and unexpected environment disturbances or instrumental anomalies, which will be the main subjects of this work. Cyclostationary properties also exist mainly due to the orbital modulations of the stochastic GW foreground from compact binaries, which are not within the scope of this work and will be left for future studies.
@@
Routine maintenance, such as antenna re-orientations, payloads calibrations and etc., could cause important and even continuous disturbances of satellite platforms and affect heavily the performances of the key payloads, therefore result into varying noise auto-correlations and even data loss or gaps. For example, the scheduled re-orientations of the transmission antennas would lead to data gaps about 3.5 hours per week or 7 hours every two-weeks [33, 37]. In addition, unforeseen anomalies or failures due to hardware problems would take place and produce random unscheduled data gaps. Based on the experiences from LPF [38, 34, 24, 36], GRACE/GFO [39, 40] and also Taiji-1 [41], instrumental transients may also contaminate the science data, such as the acceleration glitches in gravitational reference sensor (GRS) systems, phase jumps in interferometers, and etc. Early studies on data processing algorithms against such possible anomalies can be found in [33, 35, 36]. Recently, Dey et al. [23] have used Bayesian analysis techniques to assess the impact of data gaps and concluded that the effects of unscheduled gaps were often more significant, and Spadaro et al. [42] used a joint parameter estimation approach to evaluate the influences of GRS glitches demonstrating that it is possible to accurately identify glitches in the absence of overlapping GW signals. These studies treated the above non-stationarities independently, and, for instrumental transients, prior knowledge of transient models are required. In this work, given the potentials and capacities of deep neural networks, we seek to develop an anomaly-model-independent signal extraction method for space GW antennas that can thoroughly treat and overcome the impacts of the above three types of data non-stationarities or anomalies.
@@
Deep learning methods have already achieved considerable success in data processing of GW detections [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Compared to traditional algorithms, such as matched filtering [55, 56, 57], which generally require a vast template bank and thus consume substantial computational resources, deep learning methods are well known for their fast processing speed and excellent generalization ability, making it possible to achieve rapid and anomaly-model-independent signal extraction. For GW detection in space, Conv-TasNet [54] has demonstrated its performance in various scenarios, claiming an extraction speed improvement of up to times compared to traditional methods, however, robustness tests under anomalous conditions are overlooked. In this work, focusing on the three types of possible non-stationarities for LISA or LISA-like missions, including gaps, transients (glitches), and time varying noise auto-correlations, we developed a deep learning model for accurate extractions of coalescing MBHB signals confronted with such anomalous scenarios. This also provide new explorations into the robustness studies of deep learning models for data processing in space-borne GW missions. Our model exhibits the same performance as the Conv-TasNet model [54] does for the ideal and anomaly free scenario, while shows remarkable adaptability in signal extractions against all three types of non-stationarities and even their mixtures.
@@ III Data preparations
For LISA and LISA-like missions, laser frequency instability noise is dominant in the science measurements, and a post-processing technology called Time Delay Interferometry (TDI) [75, 76, 17] is employed to efficiently suppress such noise. Next important noises including clock noises, tilt-to-length noises and etc. could also be resolved within the TDI framework [77, 78], and the noise floor of the output data from certain TDI channel is in principle determined by combinations of GRS residual acceleration noises and optical readout noises . Following this, our data is generated through the first generation noise orthogonal TDI channels and , see [75] for details. The total noise Power Spectral Density (PSD) of the , channels are the combinations of the PSDs of the two noise components and , see Ref [79, 80, 81] for their specific combination forms, and according to the designs of LISA [19] and Taiji [82] we assume the nominal values
The transformed signals from incident GWs through the and channels reads
where is the transfer function combining antenna responses and TDI transformations, and is the polarized wave-train of the coalescing MBHB GW signal. The PyCBC package[83] and the waveform model SEOBNRv4 [84] are employed to generate the coalescing MBHB GW templates.
We established the parameter space for coalescing MBHBs as in Table 1, and employed a uniform grid to generate the diverse waveforms. Parameters such as times and phases at coalescence, sky locations, polarization angles, as well as inclinations are drawn from uniform distributions. The luminosity distances are used as a normalization factor to adjust the SNRs of signals.
@@
We inject the signal with specific SNR
where represents the signal template through channel , the inner product is defined as,
here Hz and Hz. The tilde symbol represents the Fourier transform and the complex conjugation. We set the SNR level for the training dataset as 50, and generate 10,000 samples. All samples are subjected to whitening processing, and a Tukey window with is adopted. The overlap between the output signal from our model and a template waveform can be evaluated in terms of the overlap function
@@
Following the same routine, we produce 2,000 test samples, see Fig 2 for illustration. In order to assess the model’s performance for signals extractions with different levels of SNRs, these samples were generated with SNRs ranging from 30 to 70.
@@ IV Training strategy
The MSE is used for the optimizations of individual data points, while the fractal Tanimoto similarity coefficient is for the optimizations of the overall data. Tanimoto coefficient is commonly used to measure the similarity between two vectors, and the fractal coefficient, as an enhanced version, introduces a parameter to facilitate deeper levels of optimizations. This approach ensures that the extracted data achieves good amplitude and phase agreements with the template data on a global level.
Another worth-noting modification is that, unlike in the original work [70] where the weight is set to , we set equal to 1 in our approach. This is because, the long observation time in space borne GW detection will give rise to large number of data points that having amplitudes close to zero in the generated templates, the high weights, such as , of these data points will consequently prevent the network from converging. For the parameter , which is used to adjust the depth of optimization in similarity, is initially set to 0. As the training converges, grows gradually to facilitate deeper levels of optimization.
@@ V.1 Data gaps
Following the discussions in [23], we use window functions to generate data gaps (see Ref [23] for the specific forms), which, as discussed in Sec. I can be classified into two types, that the scheduled prolonged gaps and unscheduled random but short gaps. For the scheduled gaps, we emulate an extreme situation, resulting into a data loss of 7 hours during the inspiral phases. For unscheduled gaps, we set each gap to cause 5 to 8 minutes of data loss, and progressively increase the frequency of occurrences of such randomly happened gaps to a threshold at which our model can hardly tolerate. Moreover, the unscheduled gaps are allowed to break out at any place of our data, while a prolonged data gap during the merger phase is ignored for obvious reason. Our extraction results are presented in Fig 4. Our model can proficiently extract the signals despite facing these two types of data gaps.
@@
Type of gap | Overlaps 0.9 |
No gap | 99% |
Scheduled gap | 99% |
Gap every 6 hours | 86% |
Gap every 5 hours | 83% |
Gap every 4 hours | 77% |
@@ V.3 Glitches
where is the time of glitch injection, is the impulse transferred by the glitch (i.e., the gain of test mass velocity caused by the glitch), are time scales, and denotes the Heaviside step function. Glitches are simulated and injected to one of the test mass interferometer measurement. The detailed transfer functions of glitches to the TDI outputs can be found in Ref. [35]. We considered two types of glitches according to Ref. [36], that the short-duration glitches lasting approximately 70 seconds and long-duration glitches for about 3 hours. To facilitate our evaluation, we have set the initial amplitude of both types of glitches to . Our model proficiently extracts the coalescing MBHB signals given such instrumental transients. Fig 9 shows an example of the model’s extraction result.
Duration | Peak value(ms) | Overlaps 0.9 |
None | None | 99% |
Short | 99% | |
Short | 99% | |
Short | 95% | |
Short | 72% | |
Long | 99% | |
Long | 71% |
The limits of our model’s performance against glitches is shown in Fig 10. The results indicate that the performance deteriorates as the duration of the glitch increases, and detailed results are presented in Tab 4. Specifically, for short-duration glitches, the maximum tolerable magnitude of the glitch is about , while for long-duration glitches, the maximum tolerable magnitude is about .
@@ V.4 Mixing three types of non-stationarities
Data gaps | Non-stationary Gaussian noise | Glitches |
Overlaps 0.9 |
Scheduled gap |
Unscheduled gap |
Duration |
Peak value (ms) |
1x |
3x |
Short |
83% |
1x |
3x |
Long |
81% |
Gap every 10 hours |
1x |
2x |
Short |
81% |
Gap every 10 hours |
1x |
2x |
Long |
81% |
@@ VI Concluding remarks
Deep learning has provided a method for the rapid detection of space-borne GW missions, and robustness research is key to ensuring that deep learning models are applicable in real detection scenarios. It is planned to process longer datasets in future works and to incorporate a greater variety of realistic simulation scenarios, including various levels of SNR and types of anomalies. Additionally, we will focus on unscheduled gaps, seeking new improvements to mitigate the substantial impact of such anomalies on model performance.

### v2

@@ Abstract
Sapce-borne gravitational wave antennas, such as LISA and LISA-like mission (Taiji and Tianqin), will offer novel perspectives for exploring our Universe while introduce new challenges, especially in data analysis. Aside from the known challenges like high parameter space dimension, superposition of large number of signals etc., gravitational wave detections in space would be more seriously affected by anomalies or non-stationarities in the science measurements. Considering the three types of foreseeable non-stationarities including data gaps, transients (glitches), and time-varying noise auto-correlations, which may come from routine maintenance or unexpected disturbances during science operations, we developed a deep learning model for accurate signal extractions confronted with such anomalous scenarios. Our model exhibits the same performance as the current state-of-the-art models do for the ideal and anomaly free scenario, while shows remarkable adaptability in extractions of coalescing massive black hole binary signal against all three types of non-stationarities and even their mixtures. This also provide new explorations into the robustness studies of deep learning models for data processing in space-borne gravitational wave missions.
@@ I INTRODUCTION
Since the first landmark event GW150914 captured by Adv-LIGO [1], today nearly a hundred Gravitational Wave (GW) signals from mergers of stellar mass compact binaries had been observed by the LIGO- Virgo collaboration, which lead to numerous accomplishments in astrophysics and fundamental physics [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. However, limited by seismic noises, ground-based GW detectors are typically sensitive to signals at frequencies higher than 10 Hz [15, 16, 17]. To enclose the exciting sources of larger and heavier astrophysical systems, one needs to explore the low frequency band of the GW spectrum with detectors of much longer baselines. The first generation space-borne antennas, including LISA (Laser Interferometer Space Antenna) [18, 19] and the LISA-like missions Taiji [20, 21] and Tianqin [22], would be launched in the 2030s and cover the mHz band. Offering novel perspectives for exploring the Universe, such space-borne antennas will also introduce new challenges [19], especially in data analysis [23, 24, 25].
Space antennas are supposed to response to a large amount of superimposed GW signals, including coalescing (super) Massive Black Hole Binaries (MBHB), extreme mass ratio inspirals, galactic compact binaries, evolving topological defects, primordial GW background, and also un-modeled sources. To achieve the expected scientific objectives, throughout analysis of the observational data and estimations of the relevant parameters from the largely superimposed GW signals against complicated noises are needed. Many works have been done to address these issues [26, 27, 28], and in present days the most accurate results can be obtained by matched filtering method for signals buried in Gaussian noises [29, 30]. However, the construction of a full-fledged pipeline of scientific data analysis for space-borne antennas remains still an unfinished task, especially with the boost, in recent years, from machine learning methods been included. Following the studies in [31, 32], accurate signal waveform extractions can be viewed as an intermediate step of the subsequent high-precision parameter estimations with machine learning methods, and the improvements in its performances and robustness are worth further explorations and investigations.
@@
As lessons learned from the LIGO- Virgo observatories [33] and precedent missions like GRACE/GFO [34], LISA PathFinder (LPF) [35] and also Taiji-1 [36], new issues in data analysis for space-borne GW detections begin to draw more attentions [37, 38, 39, 40, 25]. As mentioned, to precisely measure the related parameters and infer the physical properties of the sources, continues measurements without disruptions in the data are normally demanded, which however imposes hard challenges on the long-term stability and robustness of both the payloads and satellite platforms. GW detection in space would be more seriously affected by anomalies or non-stationarities in the science measurements, including instrument transients or noise bursts, data gaps, slowly varying noise auto-correlations, trends etc., as the detailed knowledge of instrumental noises is crucial to GW signals extraction and parameter estimations. According to the designs of the LISA and Taiji missions, the foreseeable anomalies or non-stationarities in science measurements may come from the routine maintenance of the satellites or the onboard payloads and unexpected environment disturbances or instrumental anomalies, which will be the main subjects of this work. Cyclostationary properties also exist mainly due to the orbital modulations of the stochastic GW foreground from compact binaries, which are not within the scope of this work and will be left for future studies.
@@
Routine maintenance, such as antenna re-orientations, payloads calibrations etc., could cause important and even continuous disturbances of satellite platforms and affect heavily the performances of the key payloads, therefore result into varying noise auto-correlations and even data loss or gaps. For example, the scheduled re-orientations of the transmission antennas would lead to data gaps about 3.5 hours per week or 7 hours every two-weeks [23]. In addition, unforeseen anomalies or failures due to hardware problems would take place and produce random unscheduled data gaps. Based on the experiences from LPF [41, 38, 24, 40], GRACE/GFO [42, 43] and also Taiji-1 [44], instrumental transients may also contaminate the science data, such as the acceleration glitches in gravitational reference sensor (GRS) systems, phase jumps in interferometers etc. Early studies on data processing algorithms against such possible anomalies can be found in [37, 39, 40]. Recently, Dey et al. [23] have used Bayesian analysis techniques to assess the impact of data gaps and concluded that the effects of unscheduled gaps were often more significant, and Spadaro et al. [45] used a joint parameter estimation approach to evaluate the influences of GRS glitches demonstrating that it is possible to accurately identify glitches in the absence of overlapping GW signals. These studies treated the above non-stationarities independently, and, for instrumental transients, prior knowledge of transient models are required. In this work, given the potentials and capacities of deep neural networks, we seek to develop an anomaly-model-independent signal extraction method for space GW antennas that can thoroughly treat and overcome the impacts of the above three types of data non-stationarities or anomalies.
@@
Deep learning methods have already achieved considerable success in data processing of GW detections [46, 47, 32, 48, 49, 50, 51, 52, 53, 54, 55, 31]. Compared to traditional algorithms, such as matched filtering [29, 30, 56], which generally require a vast template bank and thus consume substantial computational resources, deep learning methods are well known for their fast processing speed and excellent generalization ability, making it possible to achieve rapid and anomaly-model-independent signal extraction. For GW detection in space, Conv-TasNet [31] has demonstrated its performance in various scenarios, claiming an extraction speed improvement of up to times compared to traditional methods, such as the Markov-chain Monte Carlo and genetic algorithms [57, 58, 59, 60, 61], however, robustness tests under anomalous conditions are overlooked. In this work, focusing on the three types of possible non-stationarities for LISA or LISA-like missions, including gaps, transients (glitches), and time varying noise auto-correlations, we developed a deep learning model for accurate extractions of coalescing MBHB signals confronted with such anomalous scenarios. This also provide new explorations into the robustness studies of deep learning models for data processing in space-borne GW missions. Our model exhibits the same performance as the Conv-TasNet model [31] does for the ideal and anomaly free scenario, while shows remarkable adaptability in signal extractions against all three types of non-stationarities and even their mixtures.
@@ III Data preparations
For LISA and LISA-like missions, laser frequency instability noise is dominant in the science measurements, and a pre-processing technology called Time Delay Interferometry (TDI) [79, 80, 81] is employed to efficiently suppress such noise. Following this, our data is generated through the first generation and noise orthogonal TDI channels and , see [79] for details. The transformed signals from incident GWs through the and channels reads
where is the transfer function combining antenna responses and TDI transformations, and represents the polarization components of the coalescing MBHB GW signal. Specific details of this transfer function can be found in [81, 79]. The PyCBC package [82] and the waveform model SEOBNRv4 [83] are employed to generate the coalescing MBHB GW templates . We show in Fig. 2 how the template waveform will change when applying the TDI combinations. The noise floor of the output data from certain TDI channel is in principle determined by the corresponding TDI combinations of GRS residual acceleration noises and optical readout noises . Therefore, the total noise Power Spectral Density (PSD) of the , channels are the TDI combinations of the PSDs of the two noise components and , see Ref [81, 79, 84, 85] for their specific combination forms, and according to the designs of LISA [19] and Taiji [86] we assume the nominal values
The transformed waveform data from a certain channel will be superimposed with the noises that go through the same TDI channel, which will serve as the simulated inputs into our model.
We established the parameter space for coalescing MBHBs as in Table 1, and employed a uniform grid to generate the diverse waveforms. Parameters such as times and phases at coalescence, sky locations, polarization angles, as well as inclinations are drawn from uniform distributions.
@@
We inject the signal with specific optimal Signal-to-Noise Ratios (SNR)
where represents the signal template through the TDI channel , the inner product is defined as,
here, Hz and Hz. The tilde symbol represents the Fourier transform and the complex conjugation. We set the SNR level for the training dataset as 50, and generate 10,000 samples. The distances of the sources are adjusted according to the specific optimal SNR. All samples are subjected to whitening processing, and a Tukey window with is adopted. The overlap between the signal waveform extracted by the model from the TDI output data and the whitened template waveform through channel or can be evaluated in terms of the overlap function
@@
Following the same routine, we produce 2,000 test samples, see Fig 3 for illustration. In order to assess the model’s performance for signals extractions with relative low SNRs and test the robustness of our model against various data anomalies, these samples were generated with SNRs ranging from 30 to 70.
@@ IV Training strategy
The MSE is used for the optimizations of individual data points, while the fractal Tanimoto similarity coefficient is used for the optimizations of the overall data. Tanimoto coefficient is commonly used to measure the similarity between two vectors, and the fractal coefficient, as an enhanced version, introduces a parameter to facilitate deeper levels of optimizations. This approach ensures that the extracted data achieves good amplitude and phase agreements with the template data on a global level.
Another worth-noting modification is that, unlike in the original work [74] where the weight is set to , we set equal to 1 in our approach. This is because, the long observation time in space borne GW detection will give rise to large number of data points that having amplitudes close to zero in the generated templates, the high weights, such as , of these data points with will reduce the weights of the data from the merger phases, and therefore the optimizations will be highly affected by the data segments away from the important merge phases and consequently prevent the network from converging. For the parameter , which is used to adjust the depth of optimization in similarity, is initially set to 0. As the training converges, grows gradually to facilitate deeper levels of optimization.
@@ V.1 Data gaps
Following the discussions in [23], we use window functions to generate data gaps (see Ref [23] for the specific forms), which, as discussed in Sec. I can be classified into two types, that the scheduled prolonged gaps and unscheduled random but short gaps. For the scheduled gaps, we emulate an extreme situation, resulting into a data loss of 7 hours during the inspiral phases. For unscheduled gaps, it can last up to days or cause data loss of a few minutes per day due to glitch masking according to [37]. Considering the length of our data, here we choose the second type of unscheduled gap. We set each gap to cause 5 to 8 minutes of data loss, and progressively increase the frequency of occurrences of such randomly happened gaps to a threshold at which our model can hardly tolerate. Moreover, the unscheduled gaps are allowed to break out at any place of our data, while a prolonged data gap during the merger phase is ignored for obvious reason. Our extraction results are presented in Fig 5. Our model can proficiently extract the signals despite facing these two types of data gaps.
@@
Type of gap | Overlaps 0.9 | Mean | Std. Dev. |
No gap | 100.0% | 0.997 | 0.005 |
Scheduled gap | 99.9% | 0.996 | 0.016 |
Gap every 6 hours | 86.7% | 0.935 | 0.169 |
Gap every 5 hours | 83.1% | 0.923 | 0.181 |
Gap every 4 hours | 77.3% | 0.893 | 0.220 |
@@ V.3 Glitches
where is the time of glitch injection, is the impulse transferred by the glitch (i.e., the gain of test mass velocity caused by the glitch), are time scales, and denotes the Heaviside step function. Glitches are simulated and injected to one of the test mass interferometer measurement. The detailed transfer functions of glitches to the TDI outputs can be found in Ref. [39]. We considered two types of glitches according to Ref. [40], that the short-duration glitches lasting approximately 70 seconds and long-duration glitches for about 3 hours. The initial amplitudes of both types of glitches were set to , see Tab 4 for detailed parameters, and we tested the robustness of our model by increasing the amplitudes by 0.5 times for each new experiment until the performance fell below the threshold we set.
Duration | (m/s) | (s) | (s) |
Short | 10 | 11 | |
Long | 2000 | 2001 |
Our model proficiently extracts the coalescing MBHB signals given such instrumental transients. Fig 10 shows an example of the model’s extraction result.
Duration | Peak value(ms-2) | Overlaps 0.9 | Mean | Std. Dev. |
None | None | 100.0% | 0.997 | 0.005 |
Short | 100.0% | 0.997 | 0.005 | |
Short | 99.9% | 0.995 | 0.010 | |
Short | 94.9% | 0.977 | 0.057 | |
Short | 72.0% | 0.893 | 0.161 | |
Long | 100.0% | 0.996 | 0.006 | |
Long | 31% | 0.830 | 0.112 |
The limits of our model’s performance against glitches is shown in Fig 11. The results indicate that the performance deteriorates as the duration of the glitch increases, and detailed results are presented in Tab 5. Specifically, for short-duration glitches, the maximum tolerable magnitude of the glitch is about , while for long-duration glitches, the maximum tolerable magnitude is about . We can see that the long-duration glitch has a greater impact on the extraction results. This is because we adjust its duration to be consistent with the duration of the merge phase of the MBHB signal, which greatly affects the model’s judgment, resulting in erroneous extraction.
@@ V.4 Mixing three types of non-stationarities
Data gaps | Non-stationary Gaussian noise | Glitches | Results | |||||
Scheduled gap | Unscheduled gap | Duration | Peak value (ms-2) | Overlaps 0.9 | Mean | Std. Dev. | ||
1x | 3x | Short | 81.6% | 0.919 | 0.167 | |||
1x | 3x | Long | 80.5% | 0.921 | 0.160 | |||
Gap every 10 hours | 1x | 2x | Short | 82.1% | 0.914 | 0.194 | ||
Gap every 10 hours | 1x | 2x | Long | 81.8% | 0.905 | 0.216 |
@@ VI Concluding remarks
Deep learning has provided a method for the rapid detection of space-borne GW missions, and robustness research is key to ensuring that deep learning models are applicable in real detection scenarios. In the future, we will use this signal extraction model as a pre-processing model to reduce the difficulty of subsequent scientific analysis such as GW signal detection and parameter estimation through robust extraction of GW signals, becoming an important part of the Taiji detection pipeline. It is planned to process longer datasets in future works and to incorporate a greater variety of realistic simulation scenarios, including various levels of SNR and types of anomalies. Additionally, we will focus on unscheduled gaps, seeking new improvements to mitigate the substantial impact of such anomalies on model performance.